Write the first term and common difference. The pattern starts at 5 and add 2 each time
1
Expert's answer
2020-02-26T11:02:47-0500

As per the given question patterns starts from 5 and we have to add 2 in each term,

then the following series we obtain

5, (5+2)=7, (7+2)=9, .........

5,7,9,11,13.............

here the above series is regarded as the arithmetic progression

with a as the first term and d = a2- a1 as the common difference

so, a= first term = 5,

common difference= (7-5)= 2.

We can find any term using the formula

an= a +(n-1)d,

an=5+2(n-1),

an=3+2n,

where an= nth term of AP.